OpenStudy (anonymous):

Winnie poured 14 cups of water into a rectangular container measuring 13 inches by 7 inches by 6 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points)

OpenStudy (anonymous):

Please help

OpenStudy (wolf1728):

change 6 cm to inches 6 cm • 1 inch / 2.54 cm = 2.3622047244 inches

OpenStudy (wolf1728):

Rectangular tank measures =13*7*2.36 = 214.76 cubic inches

OpenStudy (anonymous):

are either of those the answers or do we need to do more work

OpenStudy (wolf1728):

A) max volume = 214.76 cubic inches = =214.76 / 14.44 = 14.87 cups

OpenStudy (wolf1728):

13 x 7 x 6cm It doesn't say which of those dimensions is the height so part B can't be answered.

OpenStudy (anonymous):

Sarah wants to buy the car that has a better mileage rating. Cars of Company A travel 20 miles per gallon and cars of Company B travel 14 km per liter. [1 mile = 1.6 km; 1 gallon = 3.79 liters] Part A: Which company's car should Sarah buy? Justify your answer by comparing the mileages after converting to the same units. (5 points) Part B: If 10 gallons of gas is filled in the car which gives better mileage, and you went on a drive and checked on the gas used every 30 miles, describe appropriate scales on the x-axis and y-axis to graph the rate at which the car uses gas.

OpenStudy (anonymous):

only need part b

OpenStudy (wolf1728):

14 km / liter 1/1.6 = .625 miles per km 1/ 3.79 liters = .264 liters in a gallon so 14 km / liter = 14*.625 / .264 = 33.14 miles per gallon So this is the car that gets the better mileage.

OpenStudy (wolf1728):

Here's the graph you'll need.